# Solving Fredholm Integral Equations of the Second Kind via Wasserstein Gradient Flows

### v1

@@ D.3 Proof of Proposition 6
$(1/2)\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $C_{0}(\mathcal{V}_{t}^{N})^{1/2}-\alpha\mathtt{m}\mathcal{V}_{t}^{N}+\alpha\left\|\nabla U(0)\right\|(\mathcal{V}_{t}^{N})^{1/2}+\alpha\mathtt{c}+\alpha d$
$=$ $(C_{0}+\alpha\left\|\nabla U(0)\right\|)(\mathcal{V}_{t}^{N})^{1/2}+\alpha\mathtt{c}+\alpha d-\alpha\mathtt{m}\mathcal{V}_{t}^{N}$
Noting that for any $\mathsf{a},\mathsf{b}$ such that $\mathsf{ab}\geq 1/2$, for any $x\geq 0$ we have $\sqrt{x}\leq\mathsf{a}+\mathsf{b}x$, and setting $\mathsf{a}=(C_{0}+\alpha\left\|\nabla U(0)\right\|)/\alpha\mathtt{m}$ and $\mathsf{b}=1/2\mathsf{a}$ we have:
$(1/2)\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $\frac{(C_{0}+\alpha\left\|\nabla U(0)\right\|)^{2}}{\alpha\mathtt{m}}+\alpha\mathtt{c}+\alpha d-\frac{1}{2}\alpha\mathtt{m}\mathcal{V}_{t}^{N}$
@@
$C=2\left[\frac{(C_{0}+\alpha\left\|\nabla U(0)\right\|)^{2}}{\alpha\mathtt{m}}+\alpha\mathtt{c}+\alpha d\right]/\alpha\mathtt{m}.$ (46)
@@
We now control each of these terms. Let $\varepsilon>0$, for sufficiently large $t\geq 0$ Proposition 3 ensures $\mathbf{W}_{1}(\pi_{t}(\pi^{\star}),\pi_{\alpha,\eta}^{\star})\leq\varepsilon$. Using the Lipschitz continuity of the drift established in (31) and Assumption 2 with [16, Lemma 20], there exists $k_{0}\in\mathbb{N}$ such that for any $k\geq k_{0}$ and any $t\in[0,T]$ we have $\mathbf{W}_{1}(\pi_{t}(\pi^{N_{k}}),\pi_{t}(\pi^{\star}))\leq C_{T}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})\leq\varepsilon$ since we assumed $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. Using Proposition 4, there exists $k_{1}\in\mathbb{N}$ such that for any $k\geq k_{1}$ we have that $\mathbf{W}_{1}(\pi_{t}^{N_{k}}(\pi^{N_{k}}),\pi_{t}(\pi^{N_{k}}))\leq\varepsilon$. Since $\pi^{N_{k}}$ is invariant for $(\bm{X}_{t}^{1:N_{k}})_{t\geq 0}$ we get that $\mathbf{W}_{1}(\pi^{N_{k}},\pi_{t}^{N_{k}}(\pi^{N_{k}}))=0$. Finally, there exists $k_{2}\in\mathbb{N}$ such that for any $k\geq k_{2}$ we have $\mathbf{W}_{1}(\pi^{\star},\pi^{N_{k_{2}}})\leq\varepsilon$ since we assumed $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. Combining these results in (47), we get that $\mathbf{W}_{1}(\pi^{\star},\pi_{\alpha,\eta}^{\star})\leq 4\varepsilon$. Therefore, since $\varepsilon>0$ is arbitrary, we have that $\pi^{\star}=\pi_{\alpha,\eta}^{\star}$, which concludes the proof.

### v2

@@ E.3 Proof of Proposition 7
$\frac{1}{2}\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $C_{0}(\mathcal{V}_{t}^{N})^{1/2}-(1+\alpha)\mathtt{m}\mathcal{V}_{t}^{N}+(1+\alpha)\left\|\nabla U(0)\right\|(\mathcal{V}_{t}^{N})^{1/2}+(1+\alpha)(\mathtt{c}+d)$
$=$ $(C_{0}+(1+\alpha)\left\|\nabla U(0)\right\|)(\mathcal{V}_{t}^{N})^{1/2}+(1+\alpha)(\mathtt{c}+d)-(1+\alpha)\mathtt{m}\mathcal{V}_{t}^{N}$
Noting that for any $\mathsf{a},\mathsf{b}$ such that $\mathsf{ab}\geq 1/2$, for any $x\geq 0$ we have $\sqrt{x}\leq\mathsf{a}+\mathsf{b}x$, and setting $\mathsf{a}=(C_{0}+(1+\alpha)\left\|\nabla U(0)\right\|)/((1+\alpha)\mathtt{m})$ and $\mathsf{b}=1/2\mathsf{a}$ we have:
$\frac{1}{2}\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $\frac{(C_{0}+(1+\alpha)\left\|\nabla U(0)\right\|)^{2}}{(1+\alpha)\mathtt{m}}+(1+\alpha)(\mathtt{c}+d)-\frac{1}{2}(1+\alpha)\mathtt{m}\mathcal{V}_{t}^{N}$
@@
$C=\frac{2}{(1+\alpha)\mathtt{m}}\left[\frac{(C_{0}+(1+\alpha)\left\|\nabla U(0)\right\|)^{2}}{(1+\alpha)\mathtt{m}}+(1+\alpha)(\mathtt{c}+d)\right].$ (54)
@@
We now control each of these terms. Let $\varepsilon>0$, for sufficiently large $t\geq 0$ Proposition 4 ensures $\mathbf{W}_{1}(\pi_{t}(\pi^{\star}),\pi_{\alpha,\eta}^{\star})\leq\varepsilon$. Using the Lipschitz continuity of the drift established in (41) and Assumption 3 with [16, Lemma 20], there exists $k_{0}\in\mathbb{N}$ such that for any $k\geq k_{0}$ and any $t\in[0,T]$ we have
$\mathbf{W}_{1}(\pi_{t}(\pi^{N_{k}}),\pi_{t}(\pi^{\star}))\leq C_{T}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})\leq\varepsilon$
since we assumed $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. Using Proposition 5, there exists $k_{1}\in\mathbb{N}$ such that for any $k\geq k_{1}$ we have that $\mathbf{W}_{1}(\pi_{t}^{N_{k}}(\pi^{N_{k}}),\pi_{t}(\pi^{N_{k}}))\leq\varepsilon$. Since $\pi^{N_{k}}$ is invariant for $(\bm{X}_{t}^{1:N_{k}})_{t\geq 0}$ we get that $\mathbf{W}_{1}(\pi^{N_{k}},\pi_{t}^{N_{k}}(\pi^{N_{k}}))=0$. Finally, there exists $k_{2}\in\mathbb{N}$ such that for any $k\geq k_{2}$ we have $\mathbf{W}_{1}(\pi^{\star},\pi^{N_{k_{2}}})\leq\varepsilon$ since we assumed $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. Combining these results in (55), we get that $\mathbf{W}_{1}(\pi^{\star},\pi_{\alpha,\eta}^{\star})\leq 4\varepsilon$. Therefore, since $\varepsilon>0$ is arbitrary, we have that $\pi^{\star}=\pi_{\alpha,\eta}^{\star}$, which concludes the proof.
Appendix F Additional Experiments
We empirically validate our convergence results in Proposition 47 and (14) using the simple Gaussian example in Section 5 with $\lambda=1/2,\beta=1/2$ so that we have
$\varphi(x)$ $=\frac{1}{3}\mathcal{N}\left(x;0,1\right)$
$\mathrm{k}(x,y)$ $=\mathcal{N}\left(y;xe^{-1/2},1-e^{-1}\right).$
The unique solution is $\pi(x)=\mathcal{N}\left(x;0,1\right)$. The reference measure is $\pi_{0}(x)=\pi(x)$ so that $\pi$ is the unique minimizer of $\mathcal{F}_{\alpha}^{\eta}$ and $\alpha=0.1$. We pick a small value of $\alpha$ to analyze the behavior of the algorithm with little influence from the reference measure.
We start by checking the validity of Proposition 4, 6 on this example and consider 4 initial distributions: a diffuse Gaussian $\mathcal{N}(0,2^{2})$, a concentrated Gaussian $\mathcal{N}(0,0.1^{2})$, the target distribution $\mathcal{N}(0,1^{2})$ and a Uniform over [1,1][-1,1]. We set $N=500$, $\gamma=10^{-3}$ and iterate for $n_{T}=1000$ iterations to allow for the runs with initial distributions which are far from $\pi$ to converge.
Figure 5 shows the evolution of $\mathbf{W}_{1}(\hat{\pi}_{n}^{N},\pi_{\alpha,\eta}^{\star})$, where $\hat{\pi}_{n}^{N}$ is the output of Algorithm 1 and $\pi_{\alpha,\eta}^{\star}=\pi$ is the true minimizer, along iterations. The plot shows that $\mathbf{W}_{1}$ decreases as the number of iteration increases for all initial distributions. The exponential decay predicted by Proposition 6 is particularly evident for the concentrated and uniform initial distributions.
[…]
Figure 5: Decay of $\mathbf{W}_{1}$ along iterations for 4 initial distributions: a more diffuse Gaussian $\mathcal{N}(0,2^{2})$, a more concentrated one $\mathcal{N}(0,0.1^{2})$, the target $\mathcal{N}(0,1^{2})$ and a Uniform distribution centered at 0, $\textrm{Unif}[-1,1]$.
We also check the validity of the rate in Proposition 5 and of the convergence established in Proposition 7 on this example. We set $\gamma=10^{-2}$ and $n_{T}=200$ steps. The initial distribution is a concentrated one ($\mathcal{N}(0,0.1^{2})$) and the reference measure is $\pi_{0}=\pi$ with $\alpha=0.1$. We consider number of particles $N=50,100,200,500,1000$. Figure 6 left panel shows the evolution of $\mathbf{W}_{1}$ between the solution $\pi$ and the approximation provided by Algorithm 1 at the final time $n_{T}$. As predicted by Proposition 7, $\mathbf{W}_{1}(\hat{\pi}_{n_{T}}^{N},\pi_{\alpha,\eta}^{\star})$ decays as $N\to\infty$. The right panel shows the square root of the mean squared error ($\operatorname{MSE}$) for the mean of $\pi$. We add the theoretical rate in Proposition 5 for reference. As is often the case, for small $N$ the decay is faster than $\sqrt{N}$, but as $N\to\infty$ the $\operatorname{MSE}$ decays at the predicted rate. The constant $c_{1}(T)$ is empirically estimated to be $c_{1}(T)\approx 0.075$.
$\mathbf{W}_{1}(\hat{\pi}_{n_{T}}^{N},\pi_{\alpha,\eta}^{\star})Refer to captionNN$$\operatorname{MSE}[\textrm{mean}]$
Figure 6: Left: Decay of $\mathbf{W}_{1}$ at the final time step $n_{T}$ for increasing number of particles $N$. Right: Decay of $\operatorname{MSE}^{1/2}$ for the mean of $\pi$ for increasing number of particles NN; we add the theoretical rate $1/\sqrt{N}$ for reference. The results are averaged over 10 replicates.
We repeat the same analysis for $\gamma$ to validate the rate in (14). We use the same set up as before with $N=500$ and consider $\gamma=$. Figure 7 shows the mean squared error ($\operatorname{MSE}$) for the mean of $\pi$. We add the theoretical rate in (14) for reference.
[…]
Figure 7: Behaviour of $\operatorname{MSE}$ for the mean of $\pi$ for increasing $\gamma$; we add the theoretical rate $\sqrt{\gamma}$ for reference. The results are averaged over 10 replicates.